OpenStudy (anonymous):

At the start of the basketball game, the referee tosses a ball for a jump ball. The equation that models the height of the ball is h(t) = -16t2 + 24t + 5 During what time interval is the height of the ball above 9 feet? (Round your answer to the nearest tenth of a second.)

OpenStudy (anonymous):

set equation = 9 and solve. that is write \[-16t^2+24t+5=9\] or \[-16t^2+24t-4=0\]

OpenStudy (anonymous):

easier if you divide by -4 before solving. \[4t^2-6t+1=0\]

OpenStudy (anonymous):

then use quadratic formula

OpenStudy (anonymous):

i get \[t=\frac{3\pm\sqrt{5}}{4}\]

OpenStudy (anonymous):

ignore the - and use a calculator to find \[\frac{3+\sqrt{5}}{4}\]